OpenStudy (anonymous):

The triangle shown is a right triangle. Create the equation to be used to find the missing lengths. Do not solve the equation.

OpenStudy (anonymous):

is there an image?

OpenStudy (anonymous):

In a right triangle, all the sides correspond to and equation called pythahrian thyreom

OpenStudy (anonymous):

this equation is a^2+b^2=c^2

OpenStudy (anonymous):

c^2 is the longest side of the triangle, or the hypothenuse

OpenStudy (anonymous):

so in this triangle what is the longest side?

OpenStudy (anonymous):

y?

OpenStudy (anonymous):

yep, so c = y

OpenStudy (anonymous):

and a and b are the remaining sides

OpenStudy (anonymous):

I need to find y though

OpenStudy (anonymous):

so plugging the rest of the numbers into the equation

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

you will get 3^2 + 3^2= y^2

OpenStudy (anonymous):

6^2?

OpenStudy (anonymous):

no because 3^2 = 9, so 9 + 9=18 6^2 = 36

OpenStudy (anonymous):

so the answer is 36.

OpenStudy (anonymous):

you could do 2(3^2)

OpenStudy (anonymous):

I dont have that option

OpenStudy (anonymous):

6^2 does not equal 3^2+3^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

right now we have 3^2 + 3^2=y^2

OpenStudy (anonymous):

but we need y by itself

OpenStudy (anonymous):

so take the square root of both sides

OpenStudy (anonymous):

(but don't simplify because they ask you not to

OpenStudy (anonymous):

class is almost over. can we speed this up?

OpenStudy (anonymous):

your final answer should look like y=sqrt(3^2+3^2) or y=sqrt(2(3^2))

OpenStudy (anonymous):

\[y=\sqrt{2(3^{2})}\]

OpenStudy (anonymous):

thats the answer but the program looks a bit confusing, not sure if you can even plug that in, on paper its correct if you show your teacher